\begin{thebibliography}{99}
\linespread{0}\addtolength{\itemsep}{-1.0ex}

\bibitem{Bahouri 2011} H. Bahouri, J. Chemin, R. Danchin, Fourier Analysis and Nonlinear Partial Differential Equations. Springer-Verlag, Berlin, 2011.

\bibitem{Camassa 1993} R. Camassa, D. Holm, An integrable shallow water equation with peaked solitons. Phys. Rev. Lett., 71 (1993) 1661-1664.

\bibitem{Constantin 1998} A. Constantin, J. Escher, Wave breaking for nonlinear nonlocal shallow water equations. Acta Math., 181 (1998) 229-243.

\bibitem{Constantin 2009} A. Constantin, D. Lannes, The hydrodynamical relevance of the Camassa–Holm and Degasperis-Procesi equations.
Arch. Ration. Mech. Anal., 192 (2009) 165-186.


\bibitem{Guo 2019} Z. Guo, J. Li, Z. Yin, Local well-posedness of the incompressible Euler equations in $B_{\infty, 1}^1$ and the
inviscid limit of the Navier-Stokes equations. J. Funct. Anal., 276 (2019) 2821-2830.

\bibitem{Guo 2022} Y. Guo, W. Ye, Z. Yin, Ill-posedness for the Cauchy problem of the Camassa-Holm equation in $B_{\infty, 1}^1(\mathbb{R})$. J. Differ. Equ., 327 (2022) 127-144.

\bibitem{Guo 2023} Y. Guo, The well-posedness, ill-posedness and non-uniform dependence on initial data
for the Fornberg-Whitham equation in Besov spaces. Nonlinear Anal. RWA., 70 (2023) 103791.

\bibitem{Holmes 2016} J. Holmes, Well-posedness of the Fornberg-Whitham equation on the circle. J. Differ. Equ., 260 (2016) 8530-8549.

\bibitem{Holmes 2017}  J. Holmes, R.C. Thompson, Well-posedness and continuity properties of the Fornberg-
Whitham equation in Besov spaces. J. Differ. Equ., 263 (2017) 4355-4381.

\bibitem{Korteweg 1985} D. J. Korteweg, G. de Vries, On the change of form of long waves advancing in a rectangular canal, and on a new type of long stationary waves. Philos. Mag.,  39(240) (1895) 422-443.

\bibitem{Li 2017} J. Li, Z. Yin, Well-posedness and analytic solutions of the two-component Euler-Poincar\'{e} system. Monatsh. Math., 183 (2017)
    509-537.

\bibitem{Li 2020} J. Li, Y. Yu, W. Zhu, Non-uniform dependence on initial data for the Camassa–Holm equation in Besov spaces. J.
Differ. Equ., 269 (2020)  8686-8700.

\bibitem{Li 2023} J. Li, X. Wu, Y. Yu, W. Zhu, Norm inflation and ill-posedness for the
Fornberg-Whitham equation. J. Differ. Equ., 368 (2023) 301-316.

\bibitem{Whitham 1967} G. Whitham, Variational methods and applications to water waves. Proc. R. Soc., Math.
Phys. Eng. Sci., 299 (1967) 6-25.

\bibitem{Whitham 1978} B. Fornberg, G. Whitham, A numerical and theoretical study of certain nonlinear wave
phenomena. Philos. Trans. R. Soc. Lond. A, Math. Phys. Eng. Sci., 289 (1978) 373-404.
\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amssymb,amsmath}
\usepackage{cases}
\usepackage{amsfonts}
\usepackage{color,xcolor}
\usepackage[left=2.0cm,right=2.0cm,top=2.0cm,bottom=2.0cm]{geometry}
\usepackage[colorlinks,citecolor=blue,urlcolor=blue]{hyperref}
\usepackage[utf8]{inputenc}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}{Lemma}[section]
\newtheorem{proposition}{Proposition}[section]

\newtheorem{definition}{Definition}[section]
\newtheorem{remark}{Remark}[section]

\newcommand{\bal}{\begin{align}}
\newcommand{\bbal}{\begin{align*}}
\newcommand{\beq}{\begin{equation}}
\newcommand{\eeq}{\end{equation}}
\newcommand{\bca}{\begin{cases}}
\newcommand{\eca}{\end{cases}}

\newcommand{\pa}{\partial}
\newcommand{\fr}{\frac}

\newcommand{\dd}{\mathrm{d}}

\newcommand{\R}{\mathbb{R}}

\newcommand{\les}{\lesssim}

\newcommand{\bi}{\Big}

\begin{document}
\title{Well-posedness and Continuity Properties of the Fornberg-Whitham equation in Besov space $B^1_{\infty,1}(\R)$}

\author{Guorong Qu$^{1}$,  Xing Wu$^{2,}$\footnote{E-mail: guorongqu@163.com; ny2008wx@163.com(Corresponding author); yuxiao5726@163.com} , Y. Xiao$^{2}$\\
\small $^1$ School of Tourism Data, Guilin Tourism University, Guilin 541006, China\\
\small $^2$ \small  College of Information and Management Science,
Henan Agricultural University,\\
\small Zhengzhou, Henan, 450046, China\\}

\date{}

\maketitle\noindent{\hrulefill}

{\bf Abstract:} For the Fornberg-Whitham equation, the local well-posedness in the critical Besov space $B_{p, 1}^{1+\frac{1}{p}}(\mathbb{R})$ with $1\leq p <\infty$ has been studied in \cite{Guo 2023}(Guo,  Nonlinear Anal. RWA., 2023). However, for the endpoint case $p=\infty$, whether it is locally well-posed or ill-posed in $B_{\infty, 1}^{1}(\mathbb{R})$ is still unknown. In this paper, we prove that the Fornberg-Whitham equation is well-posed in the critical Besov space $B_{\infty, 1}^{1}(\mathbb{R})$ with solutions depending
continuously on initial data, which is different from that of the Camassa-Holm equation \cite{Guo 2022}(Guo et al.,  J. Differ. Equ., 2022). In addition, we show that this dependence is
sharp by showing that the solution map is not uniformly continuous on the initial data.

{\bf Keywords:} Fornberg-Whitham equation, well-posedness, non-uniform dependence, critical Besov spaces

{\bf MSC (2020):}35Q35, 35L30
\vskip0mm\noindent{\hrulefill}

\section{Introduction}\label{sec1}
In this paper, we consider the Cauchy problem of the  Fornberg-Whitham (FW) equation
\begin{align}\label{eq1}
\begin{cases}
u_{xxt}-u_t+\frac{9}{2}u_xu_{xx}+\frac{3}{2}uu_{xxx}-\frac{3}{2}uu_x+u_x=0, &\quad (t,x)\in \mathbb{R}^+\times\mathbb{R},\\
u(0,x)=u_0(x), &\quad x\in \mathbb{R},
\end{cases}
\end{align}
which was first introduced by Whitham  and Fornberg \cite{Whitham 1967, Whitham 1978}  to study the nonlinear wave breaking phenomena. Using the Green function $G(x)=\frac{1}{2}e^{-|x|}$ and the identity $(1-\partial_{xx})^{-1}f=G*f$ for all $f\in L^2(\mathbb{R})$, we
can rewrite (\ref{eq1}) in the following non-local form
\begin{align}\label{eq2}
\begin{cases}
u_t+\frac32uu_x=\partial_x(1-\partial_{xx})^{-1}u=\partial_xG*u, &\quad (t,x)\in \R^+\times\R,\\
u(0,x)=u_0(x), &\quad x\in \R.
\end{cases}
\end{align}
In contrast to the famous Korteweg-de Vries (KdV) equation \cite{Korteweg 1985}
 \begin{align*}
   u_t+6uu_x=-\partial_{xxx}u,
 \end{align*}
 which can not describe the wave breaking phenomena, and  the classical Camassa-Holm (CH) equation \cite{Camassa 1993, Constantin 1998, Constantin 2009}
  \begin{align*}
   u_t+uu_x=-\partial_x(1-\partial_{xx})^{-1}(u^2+\frac{1}{2}u_x^2),
 \end{align*}
the FW equation not only admits solitary traveling wave solutions like
the KdV equation, but also has peaked traveling wave solutions
as the CH equation which are of the form $u(t, x)=\frac{8}{9}e^{-\frac{1}{2}|x-\frac{4}{3}t|}\cite{Whitham 1978}.$ Moreover, unlike the KdV and CH equations that are integrable and have infinitely many conserved quantities, the FW equation is not integrable. Although the FW equation is presented in a simple form,
the only useful conservation law we know so far is $\|u\|_{L^2}.$ Therefore, the analysis of the FW
equation would be somewhat more difficult.

The local posedness (existence, uniqueness and continuous Dependence) was first established by Holmes\cite{Holmes 2016} in the Sobolev spaces $H^s$ with $s>\frac{3}{2}$ on the circle, and  later extended by Holmes-Thompson \cite{Holmes 2017} to the Besov space $B_{2, r}^s$ with $s>\frac{3}{2}, 1<r<\infty$ or $s=\frac{3}{2}, r=1.$  Recently, Guo\cite{Guo 2023} investigated the  FW equation in  a general Besov space, and obtained the local well-posedness in $B_{p, r}^s(\mathbb{R})$, $s>1+\frac{1}{p}$, $1\leq p\leq \infty,$ $1\leq r< \infty$ and critical Besov spaces $B_{p, 1}^{1+\frac{1}{p}}(\mathbb{R})$ with $1\leq p <\infty.$ Li et al.\cite{Li 2023} proved that  the  FW equation  is ill-posed in Besov space $B_{p, 1}^s(\mathbb{R})$ with $1\leq s<1+\frac{1}{p}$ and $2\leq p<\infty$ due to the norm inflation. For the endpoint case $p=\infty$, whether it is locally well-posed or ill-posed in $B_{\infty, 1}^{1}(\mathbb{R})$ is unknown. Guo et al. \cite{Guo 2022} proved the ill-posedness for the CH equation in $B_{\infty, 1}^{1}(\mathbb{R})$ due to the norm inflation. However, motivated by \cite{Guo 2019}, by fully utilizing the structure of the equation and the internal properties of $u_0$ lying in  $B_{\infty, 1}^{1}(\mathbb{R})$ , we show the well-posedness for the FW equation in critical Besov spaces $B_{\infty, 1}^{1}(\mathbb{R})$.

 Our main result is stated as follows.
\begin{theorem}\label{the1.1} Let $u_0\in B_{\infty, 1}^{1}(\mathbb{R}).$ Then there exists a time $T>0$ such that  the Cauchy problem (\ref{eq2}) has a unique solution in $\mathcal{C}([0, T];B_{\infty, 1}^{1}(\mathbb{R}))\cap \mathcal{C}^1([0, T];B_{\infty, 0}^{1}(\mathbb{R}))$ and
 the data-to-solution map $u_0\rightarrow \mathcal{S}_t(u_0)$ is continuous from any bounded subset of  $B_{\infty, 1}^{1}(\mathbb{R})$ to $\mathcal{C}([0, T];B_{\infty, 1}^{1}(\mathbb{R}))$. That is, the FW equation is locally well-posed in $B_{\infty, 1}^{1}(\mathbb{R})$ in the sense of Hadamard.
\end{theorem}
\begin{remark}\label{rem1}
Since the CH equation is locally ill-posed in $B_{\infty, 1}^{1}(\mathbb{R})$, but for FW equation, we obtain the local well-posedness in $B_{\infty, 1}^{1}(\mathbb{R})$.   This interesting fact illustrates that there is a nature difference between the two equations.
\end{remark}

From our well-posedness result, we can further show that the dependence of the solutions on initial data in $B_{\infty, 1}^{1}(\mathbb{R})$ can not be better than continuous. More precisely, we shall prove the following result.
\begin{theorem}\label{the1.2} The data-to-solution map $u_0\rightarrow \mathcal{S}_t(u_0)$ for the FW equation, defined by the Cauchy
problem (\ref{eq2}),  is not uniformly continuous from any
bounded subset in  $B_{\infty, 1}^{1}(\mathbb{R})$ into $\mathcal{C}([0, T];B_{\infty, 1}^{1}(\mathbb{R}))$. More precisely,
there exist two sequences of solutions $f_n$ and $g_n$ such that
\begin{align*}
 \|f_n\|_{B_{\infty, 1}^{1}}\lesssim 1 \qquad \qquad  and  \qquad \qquad \lim_{n\rightarrow\infty}\|g_n\|_{B_{\infty, 1}^{1}}=0,
\end{align*}
but
\begin{eqnarray*}
  \liminf_{n\rightarrow \infty} \|\mathcal{S}_t(f_n+g_n)-\mathcal{S}_t(f_n)\|_{B_{\infty, 1}^1}\gtrsim t, \quad t\in [0, T_0],
        \end{eqnarray*}
with small positive time $T_0$ for $T_0\leq T$.
\end{theorem}

{\bf Notations}:  Given a Banach space $X$, we denote the norm of a function on $X$ by $\|\|_{X}$, and \begin{eqnarray*}
\|\cdot\|_{L_T^\infty(X)}=\sup_{0\leq t\leq T}\|\cdot\|_{X}.
\end{eqnarray*}
For $\mathbf{f}=(f_1, f_2,...,f_n)\in X$,
\begin{eqnarray*}
\|\mathbf{f}\|_{X}^2=\|f_1\|_{X}^2+\|f_2\|_{X}^2+...+\|f_n\|_{X}^2.
\end{eqnarray*}
The symbol
$A\lesssim B$ means that there is a uniform positive constant $C$ independent of $A$ and $B$ such that $A\leq CB$.

\section{Preliminaries}\label{sec2}
In this section, we  review the definition of Littlewood-Paley decomposition and nonhomogeneous Besov space, and then list some useful properties. For more details, the readers can refer to \cite{Bahouri 2011}.

\begin{definition} Let $\mathcal{B}:=\{\xi\in\mathbb{R}:|\xi|\leq 4/3\}$ and $\mathcal{C}:=\{\xi\in\mathbb{R}:3/4\leq|\xi|\leq 8/3\}.$
There exist two radial functions $\chi\in C_c^{\infty}(\mathcal{B})$ and $\varphi\in C_c^{\infty}(\mathcal{C})$ both taking values in $[0,1]$ such that
$$
\chi(\xi)+\sum_{j\geq0}\varphi(2^{-j}\xi)=1 \quad \forall \;  \xi\in \R.$$
For every $u\in \mathcal{S'}(\mathbb{R})$, the Littlewood-Paley dyadic blocks ${\Delta}_j$ are defined as follows
\begin{numcases}{\Delta_ju=}
0, & if $j\leq-2$;\nonumber\\
\chi(D)u=\mathcal{F}^{-1}(\chi \mathcal{F}u), & if $j=-1$;\nonumber\\
\varphi(2^{-j}D)u=\mathcal{F}^{-1}\big(\varphi(2^{-j}\cdot)\mathcal{F}u\big), & if $j\geq0$.\nonumber
\end{numcases}
The inhomogeneous low-frequency cut-off operator $S_j$ is defined by
$$
S_ju=\sum_{q=-1}^{j-1}{\Delta}_qu.
$$

\end{definition}
\begin{definition}
Let $s\in\mathbb{R}$ and $(p,r)\in[1, \infty]^2$. The nonhomogeneous Besov space $B^{s}_{p,r}(\R)$ is defined by
\begin{align*}
B^{s}_{p,r}(\R):=\Big\{f\in \mathcal{S}'(\R):\;\|f\|_{B^{s}_{p,r}(\mathbb{R})}<\infty\Big\},
\end{align*}
where
\begin{numcases}{\|f\|_{B^{s}_{p,r}(\mathbb{R})}=}
\left(\sum_{j\geq-1}2^{jrs}\|\Delta_jf\|^r_{L^p(\mathbb{R})}\right)^{\fr1r}, &if $1\leq r<\infty$,\nonumber\\
\sup_{j\geq-1}2^{js}\|\Delta_jf\|_{L^p(\mathbb{R})}, &if $r=\infty$.\nonumber
\end{numcases}
\end{definition}
\begin{lemma}(\cite{Bahouri 2011})\label{lem2.1} \;\;Let $s>0$ and $1\leq p\leq \infty$.\\
(1) $B^s_{p,1}(\R)\cap L^\infty(\R)$ is a Banach  algebra. Moreover, $B^{0}_{\infty,1}(\R)\hookrightarrow L^\infty(\R)\hookrightarrow B^{0}_{\infty,\infty}(\R)$.\\
(2) \begin{align*}
   \lim_{j\rightarrow \infty}\|S_ju-u\|_{B_{p, 1}^s}=0
    \end{align*}
(3)For any $u,v \in B^s_{p,1}(\R)\cap L^\infty(\R)$, we have
 \begin{align*}
&\|uv\|_{B^{s}_{p,1}}\leq C(\|u\|_{B^{s}_{p,1}}\|v\|_{L^\infty}+\|v\|_{B^{s}_{p,1}}\|u\|_{L^\infty}).
\end{align*}
We also have the following interpolation inequality
\begin{align*}
&\|u\|_{B^{1}_{\infty,1}}\leq C\|u\|_{B^{0}_{\infty,\infty}}^{\fr12}\|u\|_{B^{2}_{\infty,\infty}}^{\fr12}.
\end{align*}
(4) Let $m\in \mathbb{R}$ and $f$ be an $S^m-$ multiplier (i.e., $f: \mathbb{R}\rightarrow \mathbb{R}$ is smooth and satisfies that $\forall \alpha\in \mathbb{N}$, there exists a constant $\mathcal{C}_\alpha$ such that $|\partial^\alpha f(\xi)|\leq \mathcal{C}_\alpha(1+|\xi|)^{m-\alpha}$ for all $\xi \in \mathbb{R}$). Then the operator $f(D)$ is continuous from $B_{p, 1}^s(\mathbb{R})$ to $B_{p, 1}^{s-m}(\mathbb{R})$.
\end{lemma}

\begin{lemma}\label{lem2.2}(\cite{Bahouri 2011, Li 2017})
Let $1\leq p\leq \infty$, $\sigma\geq 1+\frac{1}{p}.$
There exists a constant $C=C(p,\sigma)$ such that for any solution to the
following linear transport equation:
\begin{equation*}
\partial_t f+v\partial_x f=g,\qquad
f|_{t=0} =f_0,
\end{equation*}
the following statements hold:
\begin{align*}
  \|f(t)\|_{B^{\sigma}_{p,1}}\leq \|f_0\|_{B^{\sigma}_{p,1}}+\int_0^t\|g(\tau)\|_{B^{\sigma}_{p,1}}d\tau+\int_0^tCV^{'}(\tau)\|f(\tau)\|_{B^{\sigma}_{p,1}}d\tau
\end{align*}
or
\begin{align*}
\sup_{s\in [0,t]}\|f(s)\|_{B^{\sigma}_{p,1}}\leq e^{CV(t)}\Big(\|f_0\|_{B^\sigma_{p,1}}
+\int^t_0e^{-CV(\tau)}\|g(\tau)\|_{B^{\sigma}_{p,1}}\dd \tau\Big),
\end{align*}

with
\begin{align*}
V(t)=
\int_0^t \|\partial_x v(s)\|_{B^{\sigma-1}_{p,1}}\dd s
\end{align*}
\end{lemma}
\section{Proof of Theorem \ref{the1.1}}
\setcounter{equation}{0}

In this section, we divide Proof of Theorem \ref{the1.1} into three steps.
\subsection{Existence and Uniqueness}
To prove the local existence, it is sufficient to establish  a priori estimate of the  solution in $B_{\infty, 1}^1$.

Using Lemma \ref{lem2.2}-Lemma \ref{lem2.1}, we have
\begin{align}\label{eq3.1}
\|u(t)\|_{B^{1}_{\infty,1}}&\leq e^{CV(t)}\Big(\|u_0\|_{B^{1}_{\infty,1}}+\int_0^te^{-CV(\tau)}\|\partial_x(1-\partial_{xx})^{-1}u\|_{B^{1}_{\infty,1}}\dd \tau\Big)\nonumber\\
&\leq e^{CV(t)}\Big(\|u_0\|_{B^{1}_{\infty,1}}+\int_0^te^{-V(\tau)}\|u(\tau)\|_{B^{1}_{\infty,1}}\dd \tau\Big),
\end{align}
where $V(t)=\int_0^t \|u(\tau)\|_{B^{1}_{\infty,1}}\dd\tau$. Then we obtain from \eqref{eq3.1} that
\bbal
F(t):=e^{-CV(t)}\|u(t)\|_{B^{1}_{\infty,1}}&\leq \|u_0\|_{B^{1}_{\infty,1}}+\int_0^tF(\tau)\dd \tau,
\end{align*}
which alongs with Gronwall's inequality yields
\begin{align}\label{eq3.2}
\|u(t)\|_{B^{1}_{\infty,1}}&\leq C\|u_0\|_{B^{1}_{\infty,1}}\exp\Big(C\int_0^t \|u(\tau)\|_{B^{1}_{\infty,1}}\dd\tau\Big).
\end{align}
Let
\begin{align*}
 A(t)=C\|u_0\|_{B^{1}_{\infty,1}}\exp\Big(C\int_0^t \|u(\tau)\|_{B^{1}_{\infty,1}}\dd\tau\Big)\;\; \mbox{with}\; \;A(0)=C\|u_0\|_{B^{1}_{\infty,1}},
\end{align*}
then from (\ref{eq3.2}), we have
\begin{align*}
\frac{\dd A(t)}{\dd t}\leq CA^2(t).
\end{align*}

Solving the above differential inequalities and combining with  \ref{eq3.2}, one has
\begin{align}\label{eq3.3}
\|u(t)\|_{B^{1}_{\infty,1}}&\leq \frac{C\|u_0\|_{B^{1}_{\infty,1}}}{1-Ct\|u_0\|_{B^{1}_{\infty,1}}}
\end{align}
Fix a $T > 0$ such that $T<\frac{1}{2C\|u_0\|_{B^{1}_{\infty,1}}}$, then we have
\begin{align}\label{eq3.4}
 \|u\|_{L^\infty_TB^{1}_{\infty,1}}&\leq C\|u_0\|_{B^{1}_{\infty,1}}, \qquad \forall t\in [0, T].
\end{align}
Therefore, the solution $u$ is  uniformly bounded in $L^\infty([0, T]; B^{1}_{\infty,1}).$

The existence follows the standard procedure, we omit the details. The uniqueness is the direct result of the following lemma. In fact, suppose that $u_1,u_2\in \mathcal{C}([0,T],B^{1}_{\infty,1})$ are two solutions of (\ref{eq2}) with the same initial data $u_0$, then
we have
$$\|u_1-u_2\|_{L^\infty}\leq C\|u_1(0)-u_2(0)\|_{L^\infty}=0,$$
which implies the uniqueness.
\begin{lemma}\label{ley1} Let $u,v\in \mathcal{C}([0,T],B^{1}_{\infty,1})$ be two solutions of (\ref{eq2}) associated with $u_0$ and $v_0$, respectively. Then we have the estimate for the difference $w=u-v$
\begin{align}
&\|w\|_{L^\infty}\leq C\|w_0\|_{L^\infty},\label{eq3.5}
\\&\|w\|_{B^1_{\infty,1}}\leq C\Big(\|w_0\|_{B^1_{\infty,1}}+\int^t_0\|v_x\|_{B^1_{\infty,1}}\|w\|_{L^\infty}\dd \tau\Big),\label{eq3.6}
\end{align}
where the constants $C$ depends on $T$ and initial norm $\|u_0,v_0\|_{B^1_{\infty,1}}$.
\end{lemma}
{\bf Proof.}\quad
It is easy to check that $w$ satisfys
\begin{align}\label{eq3.7}
\begin{cases}
\pa_tw+\frac32uw_x=-\frac32wv_x+\pa_x(1-\pa_{xx})^{-1}w, \\
w(0,x)=u_0(x)-v_0(x).
\end{cases}
\end{align}
Let $p\geq 2$, taking the inner product of (\ref{eq3.7}) with $p|w|^{p-2}w$, then one has
\begin{align*}
  \frac{\dd\|w\|^p_{L^p}}{\dd t}&=\frac{3}{2}\int u_x|w|^p\dd t-\frac{3}{2}\int p|w|^pv_x\dd t+\int_t(\partial_xG*w)p|w|^{p-2}w\dd t.
\end{align*}
Using the Young inequality and H\"{o}lder inequality, we obtain
\begin{align*}
  \frac{\dd\|w\|^p_{L^p}}{\dd t}&\leq C(\|u_x\|_{L^\infty}+p\|v_x\|_{L^\infty}+p)\|w\|^p_{L^p},
\end{align*}
thus
\begin{align*}
  \frac{\dd\|w\|_{L^p}}{\dd t}&\leq C(\|u_x\|_{L^\infty}+\|v_x\|_{L^\infty}+1)\|w\|_{L^p}.
\end{align*}
Using the embedding $B_{\infty, 1}^1(\mathbb{R})\hookrightarrow \mathcal{C}^{0, 1}(\mathbb{R})$, and integrating the above differential inequality with respect to time $t\in [0, T]$ yields
\begin{align*}
\|w\|_{L^p}
&\leq \|w_0\|_{L^p}\exp\Big(C\int_0^T\big(\|v\|_{B^1_{\infty,1}}+\|v\|_{B^1_{\infty,1}}+1\big)\dd s\Big)\leq C\|w_0\|_{L^p}.
\end{align*}
Let $p\rightarrow\infty$, we obtain the desired estimate (\ref{eq3.5}).

Applying Lemma Lemma \ref{lem2.2}-Lemma \ref{lem2.1} to  (\ref{eq2}) yields
\begin{align*}
\|w(t)\|_{B^1_{\infty,1}}-\|w_0\|_{B^1_{\infty,1}}&\les\int_0^t\Big( \|u\|_{B^1_{\infty,1}}\|w\|_{B^1_{\infty,1}}+\|wv_x,\pa_x(1-\pa_{xx})^{-1}w\|_{B^1_{\infty,1}}\Big)\mathrm{d}\tau\\
&\les \int_0^t(\|u\|_{B^1_{\infty,1}}+1)\|w\|_{B^1_{\infty,1}}+\|wv_x\|_{B^1_{\infty,1}}\mathrm{d}\tau\\
&\les\int_0^t\|w\|_{B^1_{\infty,1}}\|v_x\|_{L^\infty}+\|v_x\|_{B^1_{\infty,1}}\|w\|_{L^\infty}+(\|u\|_{B^1_{\infty,1}}+1)\|w\|_{B^1_{\infty,1}}\mathrm{d}\tau\\
&\les\int_0^t\|w\|_{B^1_{\infty,1}}\big(1+\|u, v\|_{B^1_{\infty,1}}\big)\mathrm{d}\tau+\int_0^t\|v_x\|_{B^1_{\infty,1}}\|w\|_{L^\infty}\mathrm{d}\tau.
\end{align*}
Gronwall's inequality yields  the desired (\ref{eq3.6}).

\subsection{Continuous Dependence}
Now we will prove that the solution of the FW equation in $B_{\infty, 1}^1(\mathbb{R})$ is continuously dependent on the initial data by \cite{Guo 2019}. The main difficulty lies in that the FW equation is of hyperbolic type. Precisely speaking, if $u,v\in \mathcal{C}([0,T],B^{1}_{\infty,1})$ are two solutions of \eqref{eq1} associated with $u_0$ and $v_0$, in view of \eqref{eq3.6}, we have to tackle with the term $\|v_x\|_{B^1_{\infty,1}}$. To bypass this, we can take $v=\mathcal{S}_{t}(S_Nu_0)$ as the solution to \eqref{eq1} with initial data $S_Nu_0$.

Letting $u=\mathcal{S}_{t}(u_0)$ and $v=\mathcal{S}_{t}(S_Nu_0)$, using Lemma \ref{lem2.2} and \eqref{eq3.5}, we have
\bbal
&\mathcal{S}_{t}(S_Nu_0)\|_{B^2_{\infty,1}}\leq C\|S_Nu_0\|_{B^2_{\infty,1}}\leq C2^N\|u_0\|_{B^1_{\infty,1}}
\end{align*}
and
\bbal
\|\mathcal{S}_{t}(S_Nu_0)-\mathcal{S}_{t}(u_0)\|_{L^\infty}&\leq C\|S_Nu_0-u_0\|_{L^\infty}
\\&\leq C\|S_Nu_0-u_0\|_{B^0_{\infty,1}}\\
&\leq C2^{-N}\|S_{N}u_0-u_0\|_{B^1_{\infty,1}},
\end{align*}
which combining with (\ref{eq3.6}) imply
\bbal
\|\mathcal{S}_{t}(S_Nu_0)-\mathcal{S}_{t}(u_0)\|_{B^1_{\infty,1}}&\leq C\|S_Nu_0-u_0\|_{B^1_{\infty,1}}.
\end{align*}
Then, for $u_0,\widetilde{u}_0\in B^1_{\infty,1}$, we have
\bbal
\|\mathcal{S}_{t}(u_0)-\mathcal{S}_{t}(\widetilde{u}_0)\|_{B^1_{\infty,1}}
&\leq \|\mathcal{S}_{t}(S_Nu_0)-\mathcal{S}_{t}(u_0)\|_{B^1_{\infty,1}}+\|\mathcal{S}_{t}(S_N\widetilde{u}_0)-\mathcal{S}_{t}(\widetilde{u}_0)\|_{B^1_{\infty,1}}\\
&~~+\|\mathcal{S}_{t}(S_Nu_0)-\mathcal{S}_{t}(S_N\widetilde{u}_0)\|_{B^1_{\infty,1}}
\\&\leq C\|S_Nu_0-u_0\|_{B^1_{\infty,1}}+C\|S_N\widetilde{u}_0-\widetilde{u}_0\|_{B^1_{\infty,1}}\\
&~~+C\|\mathcal{S}_{t}(S_Nu_0)-\mathcal{S}_{t}(S_N\widetilde{u}_0)\|_{B^1_{\infty,1}}
\\
&:=\mathcal{I}_1+\mathcal{I}_2+\mathcal{I}_3.
\end{align*}
Using the interpolation inequality in Lemma \ref{lem2.1}, we obtain
\bbal
\mathcal{I}_3&\leq C\|\mathcal{S}_{t}(S_Nu_0)-\mathcal{S}_{t}(S_N\widetilde{u}_0)\|^{\fr12}_{B^0_{\infty,\infty}}
\|\mathcal{S}_{t}(S_Nu_0)-\mathcal{S}_{t}(S_N\widetilde{u}_0)\|^{\fr12}_{B^2_{\infty,\infty}}\\
&\leq C\|\mathcal{S}_{t}(S_Nu_0)-\mathcal{S}_{t}(S_N\widetilde{u}_0)\|^{\fr12}_{L^\infty}
\|\mathcal{S}_{t}(S_Nu_0)-\mathcal{S}_{t}(S_N\widetilde{u}_0)\|^{\fr12}_{B^2_{\infty,1}}\\
&\leq C\|S_Nu_0-S_N\widetilde{u}_0\|^{\fr12}_{L^\infty}
\|\mathcal{S}_{t}(S_Nu_0)-\mathcal{S}_{t}(S_N\widetilde{u}_0)\|^{\fr12}_{B^2_{\infty,1}}\\
&\leq C2^{\frac N2}\|u_0-\widetilde{u}_0\|^{\frac12}_{B^1_{\infty,1}},
\end{align*}
which clearly implies
\bbal
\|\mathcal{S}_{t}(u_0)-\mathcal{S}_{t}
(\widetilde{u}_0)\|_{B^1_{\infty,1}}
&\lesssim \|S_Nu_0-u_0\|_{B^1_{\infty,1}}+
\|S_N\widetilde{u}_0-\widetilde{u}_0\|_{B^1_{\infty,1}}+2^{\frac N2}\|u_0-\widetilde{u}_0\|^{\frac12}_{B^1_{\infty,1}}
\\&\lesssim \|S_Nu_0-u_0\|_{B^1_{\infty,1}}+
\|S_N\widetilde{u}_0-\widetilde{u}_0\|_{B^1_{\infty,1}}+2^{\frac N2}\|u_0-\widetilde{u}_0\|^{\frac12}_{B^1_{\infty,1}}.
\end{align*}
Taking $N$ large enough, we can conclude that the   the data-to-solution map $u_0\rightarrow \mathcal{S}_t(u_0)$ is  continuously dependent on initial data.

\section{Proof of Theorem \ref{the1.2}}
\setcounter{equation}{0}

 Let $\hat{\phi}\in \mathcal{C}^\infty_0(\mathbb{R})$ be an even, real-valued and non-negative function on $\R$ and satisfy
\begin{numcases}{\hat{\phi}(x)=}
1, &if $|x|\leq \frac{1}{4}$,\nonumber\\
0, &if $|x|\geq \frac{1}{2}$.\nonumber
\end{numcases}
\begin{lemma}\label{lem4.1}
We define the high frequency function $f_n$ and the low frequency functions $g_n$ as follows
\bbal
&f_n=2^{-n}\phi(x)\sin \bi(\frac{17}{12}2^nx\bi),\\
&g_n=\frac{12}{17}2^{-n}\phi(x),\quad n\gg1.
\end{align*}
Then for any $\sigma\in\R$, we have
\bal
&\|f_n\|_{L^\infty}\leq C2^{-n}\phi(0)\quad\text{and}\quad\|g_n\|_{L^\infty}\leq C2^{-n}\phi(0),\label{y0}\\
&\|f_n\|_{B^{\sigma}_{\infty,1}}\leq C2^{(\sigma-1)n}\phi(0)\quad\text{and}\quad\|g_n\|_{B^\sigma_{\infty,1}}\leq C2^{-(n+\sigma)}\phi(0),\label{yz}\\
&\liminf_{n\rightarrow \infty}\|g_n\pa_xf_n\|_{B^{1}_{\infty,\infty}}\geq M_1,\label{yz3}
\end{align}
for some positive constants $C, M_1$.
\end{lemma}
{\bf Proof.}\quad We refer to see Lemma 3.2-Lemma 3.4 in \cite{Li 2020} for the proof with minor modifications.
\begin{proposition}\label{pro1}
Assume that $\|u_0\|_{B^{1}_{\infty,1}}\lesssim 1$. Under the assumptions of Theorem \ref{the1.1}, we have
\bal\label{et0}
\|\mathcal{S}_{t}(u_0)-u_0-t\mathbf{v}_0(u_0)\|_{B^{1}_{\infty,1}}\leq Ct^{2}\mathcal{E}(u_0),
\end{align}
here $\mathbf{v}_0(u_0):=\partial_x(1-\partial_{xx})^{-1}u_0-\frac{3}{2}u_0\pa_x u_0$ and
\bbal
\mathcal{E}(u_0)&:=1+\|u_0\|_{L^\infty}\big(\|u_0\|_{B^{2}_{\infty,1}}+
\|u_0\|_{L^\infty}
\|u_0\|_{B^{3}_{\infty,1}}\big).
\end{align*}
\end{proposition}
{\bf Proof.}\quad For simplicity, we denote $u(t)=\mathbf{S}_t(u_0)$.
By the Mean Value Theorem and (\ref{eq3.4}), we obtain
\bal\label{et1}
\|u(t)-u_0\|_{L^\infty}&\leq\int^t_0\|\pa_\tau u\|_{L^\infty} \dd\tau
\nonumber\\&\leq \int^t_0\frac{3}{2}\|u\pa_xu\|_{L^\infty}\dd \tau+\int^t_0\|\partial_xG*u\|_{L^\infty}\dd \tau
\nonumber\\&\leq C\int^t_0\|u(\tau)\|_{L^\infty}\big(\|u(\tau)\|_{C^{0,1}}+1\big)\dd \tau\nonumber\\
&\leq Ct\|u_0\|_{L^\infty}.
\end{align}

Using Lemma \ref{lem2.1}, \eqref{eq3.4} yield
\bal\label{et2}
\|u(t)-u_0\|_{B^{1}_{\infty,1}}
&\leq \int^t_0\|\pa_\tau u\|_{B^{1}_{\infty,1}} \dd\tau
\nonumber\\&\leq \int^t_0\|\partial_x(1-\partial_{xx})^{-1}u\|_{B^{1}_{\infty,1}} \dd\tau+ \int^t_0\frac{3}{2}\|u \pa_xu\|_{B^{1}_{\infty,1}} \dd\tau
\nonumber\\&\leq Ct\big(\|u\|_{L^\infty _T(B^{1}_{\infty,1})}+\|u\|^2_{L^\infty _T(B^{1}_{\infty,1})}+\|u\|_{L^\infty _T(L^\infty)}\|u_x\|_{L_T^\infty(B^{1}_{\infty, 1})}\big)
\nonumber\\&\leq Ct\big(1+\|u_0\|_{L^\infty}
\|u_0\|_{B^{2}_{\infty,1}}\big).
\end{align}
Similarly, we have
\bal\label{et3}
\|u(t)-u_0\|_{B^{2}_{\infty,1}}
&\leq \int^t_0\|\pa_\tau u\|_{B^{2}_{\infty,1}} \dd\tau
\nonumber\\&\leq \int^t_0\|\partial_x(1-\partial_{xx})^{-1}u\|_{B^{2}_{\infty,1}} \dd\tau+ \int^t_0\frac{3}{2}\|u \pa_xu\|_{B^{2}_{\infty,1}} \dd\tau
\nonumber\\&\leq Ct\big(\|u\|_{L^\infty_T(B^{2}_{\infty,1})}
+\|u\|_{L^\infty_T(L^\infty)}\|u\|_{L^\infty_T(B^{3}_{\infty,1})}\big)
\nonumber\\&\leq Ct\big(\|u_0\|_{B_{\infty, 1}^2}
+\|u_0\|_{L^\infty}\|u_0\|_{B^{3}_{\infty,1}}\big).
\end{align}
Using the Mean Value Theorem and Lemma \ref{lem2.1} once again, we obtain that
\bal\label{et4}
\|u(t)-u_0-t\mathbf{v}_0(u_0)\|_{B^{1}_{\infty,1}}
&\leq \int^t_0\|\pa_\tau u-\mathbf{v}_0(u_0)\|_{B^{1}_{\infty,1}} \dd\tau
\nonumber\\&\leq \int^t_0\|\partial_x(1-\partial_{xx})^{-1}(u-u_0)\|_{B^{1}_{\infty,1}} \dd\tau+\int^t_0\frac{3}{2}\|u\pa_xu-u_0\pa_xu_0\|_{B^{1}_{\infty,1}} \dd\tau
\nonumber\\&\lesssim \int^t_0\|u(\tau)-u_0\|_{B^{1}_{\infty,1}} \dd\tau+\int^t_0\|u(\tau)-u_0\|_{L^\infty} \|u(\tau)\|_{B^{2}_{\infty,1}} \dd\tau
\nonumber\\&\quad \ + \int^t_0\|u(\tau)-u_0\|_{B^{2}_{\infty,1}}  \|u_0\|_{L^\infty}\dd \tau\nonumber\\
&\lesssim \int^t_0\|u(\tau)-u_0\|_{B^{1}_{\infty,1}} \dd\tau+\|u_0\|_{B^{2}_{\infty,1}}\int^t_0\|u(\tau)-u_0\|_{L^\infty}  \dd\tau
\nonumber\\&\quad \ + \|u_0\|_{L^\infty}\int^t_0\|u(\tau)-u_0\|_{B^{2}_{\infty,1}}  \dd \tau.
\end{align}
Plugging \eqref{et1}--\eqref{et3} into \eqref{et4} yields the desired result \eqref{et0}. Thus, we complete the proof of Proposition \ref{pro1}.

Now we prove the non-uniform continuous dependence.

Set $u^n_0=f_n+g_n$ and compare the solution $\mathcal{S}_{t}(u^n_0)$ with $\mathcal{S}_{t}(f_n)$. Obviously,
\bbal
\|u^n_0-f_n\|_{B^{1}_{\infty,1}}=\|g_n\|_{B^{1}_{\infty,1}}\leq C2^{-n},
\end{align*}
which means that
\bbal
\lim_{n\to\infty}\|u^n_0-f_n\|_{B^{1}_{\infty,1}}=0.
\end{align*}
From Lemma \ref{lem4.1}, one has
\bbal
&\|u^n_0,f_n\|_{B^{\sigma}_{\infty,1}}\leq C2^{(\sigma-1)n}\quad \text{for}\quad \sigma\geq1,\\
&\|u^n_0,f_n\|_{L^\infty}\leq C2^{-n},
\end{align*}
which implies
\bbal
\mathcal{E}(u^n_0)+\mathcal{E}(f_n)\leq C.
\end{align*}
Using the facts
\bbal
&\big\|u^n_{0}\pa_xg_n\big\|_{B^{1}_{\infty,1}}\leq C\big\|u^n_0\big\|_{B^{1}_{\infty,1}}\big\|g_n\big\|_{B^{2}_{\infty,1}}\leq C2^{-n},\\
&\big\|\partial_x(1-\partial_{xx})^{-1}(u_0^n-f_n)\big\|_{B^{1}_{\infty,1}}\leq C\big\|g_n\big\|_{B^{1}_{\infty,1}}\leq C2^{-n}.
\end{align*}
we deduce that
\bal\label{yyh}
\big\|\mathcal{S}_{t}(u^n_0)-\mathcal{S}_{t}(f_n)\big\|_{B^{1}_{\infty,1}}\geq&~ t\big\|g_n\pa_xf_n\big\|_{B^{1}_{\infty,1}}-t\big\|u^n_{0}\pa_xg_n,\;\partial_x(1-\partial_{xx})^{-1}(u_0^n-f_n)\big\|_{B^{1}_{\infty,1}}-Ct^{2}-C2^{-n}\nonumber\\
\geq&~ t\big\|g_n\pa_xf_n\big\|_{B^{1}_{\infty,1}}-Ct2^{-n}-Ct^{2}-C2^{-n},
\end{align}
Notice that \eqref{yz3}
\begin{eqnarray*}
      \liminf_{n\rightarrow \infty} \big\|g_n\pa_xf_n\big\|_{B^{1}_{\infty,1}}\gtrsim M_1,
        \end{eqnarray*}
then we deduce from \eqref{yyh} that
\bbal
\liminf_{n\rightarrow \infty}\big\|\mathcal{S}_t(f_n+g_n)-\mathcal{S}_t(f_n)\big\|_{B^{1}_{\infty,1}}\gtrsim t\quad\text{for} \ t \ \text{small enough}.
\end{align*}
This completes the proof of Theorem \ref{the1.2}.

\section*{Data Availability} Data sharing is not applicable to this article as no new data were created or analyzed in this study.

\section*{Conflict of interest}
The authors declare that they have no conflict of interest.

\section*{Acknowledgments}
 Y. Xiao is supported by the National Natural Science Foundation of China under Grant 11901167.

\end{document}